OpenStudy (lovelycharm):

Given the equation −4Square root of x minus 3 = 12, solve for x and identify if it is an extraneous solution.

OpenStudy (lovelycharm):

\[\sqrt[-4]{x-3}\]

OpenStudy (anonymous):

ok, let's rewrite the question into (x-3)^(-1/4) = 12 Since the exponent is negative flip the inside (1/(x-3))^(1/4) = 12 Let's take out the (1/4) by putting the power of 4 on both sides ((1/(x-3))^(1/4))^4 = 12^4 The exponents cancel out on the left side So, 1/(x-3) = 12^4 Send the x-3 to the other side and 12^4 to left side 1/12^4 = x-3 thus, x = (1/12^4)+3 The solution is not extraneous solution because there is only one value for x.

OpenStudy (lovelycharm):

wait is x = to 0 or 12 @Darkmirror

OpenStudy (anonymous):

neither. x = (1/12^4)+ 3

OpenStudy (lovelycharm):

c x = 0, solution is not extraneous x = 0, solution is extraneous x = 12, solution is not extraneous x = 12, solution is extraneous

OpenStudy (lovelycharm):

@Darkmirror

OpenStudy (anonymous):

Hmm... can you please recheck your equation? @lovelycharm

OpenStudy (lovelycharm):

u right i forget to put =12 at the end sorry

OpenStudy (lovelycharm):

@Darkmirror

OpenStudy (lovelycharm):

\[\sqrt[-4]{x-3=12}\]

OpenStudy (anonymous):

You are miswriting it. it should be -4 * sqrt(x-3) = 12 so, it's x = 12, solution is not extraneous

OpenStudy (anonymous):

Srry, x=12 solution is extraneous because it doesn't equal to 12 when you plug back in

OpenStudy (lovelycharm):

it fine thank u